# FedBiP: Heterogeneous One-Shot Federated Learning with Personalized Latent Diffusion Models

### v1

@@ Abstract
One-Shot Federated Learning (OSFL), a special decentralized machine learning paradigm, has recently gained significant attention. OSFL requires only a single round of client data or model upload, which reduces communication costs and mitigates privacy threats compared to traditional FL. Despite these promising prospects, existing methods face challenges due to client data heterogeneity and limited data quantity when applied to real-world OSFL systems. Recently, Latent Diffusion Models (LDM) have shown remarkable advancements in synthesizing high-quality images through pretraining on large-scale datasets, thereby presenting a potential solution to overcome these issues. However, directly applying pretrained LDM to heterogeneous OSFL results in significant distribution shifts in synthetic data, leading to performance degradation in classification models trained on such data. This issue is particularly pronounced in rare domains, such as medical imaging, which are underrepresented in LDM’s pretraining data. To address this challenge, we propose Federated Bi-Level Personalization (FedBiP), which personalizes the pretrained LDM at both instance-level and concept-level. Hereby, FedBiP synthesizes images following the client’s local data distribution without compromising the privacy regulations. FedBiP is also the first approach to simultaneously address feature space heterogeneity and client data scarcity in OSFL. Our method is validated through extensive experiments on three OSFL benchmarks with feature space heterogeneity, as well as on challenging medical and satellite image datasets with label heterogeneity. The results demonstrate the effectiveness of FedBiP, which substantially outperforms other OSFL methods.
@@ 1 Introduction
Federated Learning (FL) (McMahan et al., 2017) is a decentralized machine learning paradigm, in which multiple clients collaboratively train neural networks without centralizing their local data. However, traditional FL frameworks require frequent communication between a server and clients to transmit model weights, which would lead to significant communication overheads (Kairouz et al., 2021). Additionally, such frequent communication increases system susceptibility to privacy threats, as transmitted data can be intercepted by attackers who may then execute membership inference attacks (Lyu et al., 2020). In contrast, a special variant of FL, One-Shot Federated Learning (OSFL) (Guha et al., 2019), serves as a promising solution. OSFL requires only single-round server-client communication, thereby enhancing communication efficiency and significantly reducing the risk of interception by malicious attackers. Therefore, we focus on OSFL given its promising properties.
Despite these promising prospects, existing methods for OSFL face significant challenges when applied to real-world scenarios. Previous works (Guha et al., 2019; Li et al., 2020) require additional public datasets, presenting challenges in privacy-critical domains such as medical data (Liu et al., 2021), where acquiring data that conforms to client-specific distributions is often impractical. Alternatively, they can involve the transmission of entire model weights (Zhang et al., 2022) or local training data (Zhou et al., 2020), which are inefficient and increase the risk of privacy leakage. Moreover, these approaches overlook the issue of feature space heterogeneity, wherein the data features across different clients exhibit non-identically distributed properties. This presents an important and prevalent challenge as emphasized in (Li et al., 2021; Chen et al., 2023). Another vital challenge in (One-Shot) FL is the limited quantity of data available from clients (McMahan et al., 2017). This problem is particularly notable in specialized domains, such as medical or satellite imaging (So et al., 2022) where data collection is time-consuming and costly.
Data augmentation constitutes a promising strategy to address these challenges in traditional FL (Zhu et al., 2021; Li et al., 2022) by optimizing an auxiliary generative model. However, its reliance on multiple communication rounds makes it unsuitable for OSFL. Recently, diffusion models (Ho et al., 2020), particularly Latent Diffusion Model (LDM) (Rombach et al., 2022), have gained significant attention due to their capability to synthesize high-quality images after being pretrained on large-scale datasets. They are pro-
@@
Figure 1: Feature map visualization of original client images (real), synthetic images by prompted pretrained LDM (pretrained), and our method (FedBiP) on two datasets. FedBiP effectively mitigates the strong distribution shifts between pretrained LDM and client local data.
@@
ven effective in various tasks, including training data augmentation (Yuan et al., 2023; Azizi et al., 2023) and addressing feature shift problems (Niemeijer et al., 2024; Gong et al., 2023) under centralized settings. However, directly applying a pretrained LDM for specialized domains presents challenges. As demonstrated in Figure 1, there is a noticeable feature distributional shift and visual discrepancy between real and synthetic data. This mismatch could lead to performance degradation when incorporating such synthetic data into the training process, especially in heterogeneous OSFL settings, where each client possesses data with varying distributions.
@@
Therefore, in this paper, we propose Federated Bi-Level Personalization (FedBiP), a framework designed to adapt pretrained LDM for synthesizing high-quality training data that adheres to client-specific data distributions in OSFL. FedBiP incorporates personalization of the pretrained LDM at both instance and concept levels. Specifically, instance-level personalization focuses on adapting the pretrained LDM to generate high-fidelity samples that closely align with each client’s local data while preserving data privacy. Concurrently, concept-level personalization integrates category and domain-specific concepts from different clients to enhance data generation diversity at the central server. This bi-level personalization approach improves the performance of classification models trained on the synthesized data. Our contributions can be summarized as follows:
@@ 2.1 One-Shot Federated Learning
A variety of efforts have been made to address One-Shot Federated Learning (OSFL), primarily from two complementary perspectives: one focuses on model aggregation through techniques such as model prediction averaging (Guha et al., 2019), majority voting (Li et al., 2020), conformal prediction method (Humbert et al., 2023), loss surface adaptation (Su et al., 2023), or Bayesian methods (Yurochkin et al., 2019; Chen & Chao, 2020; Hasan et al., 2024). These approaches may not fully exploit the underlying knowledge across different client data distributions. Another aims to transmit training data instead of model weights: data distribution (Kasturi et al., 2020; Beitollahi et al., 2024; Shin et al., 2020), Generative Adversarial Networks (GANs) (Goodfellow et al., 2020; Zhang et al., 2022; Kasturi & Hota, 2023; Kang et al., 2023; Dai et al., 2024), or distilled dataset (Zhou et al., 2020; Song et al., 2023) are optimized and transmitted to the central server for subsequent model training. Given the success of diffusion models (Rombach et al., 2022), (Zhang et al., 2023; Yang et al., 2024b) suggests transmitting image captions to reproduce training data at the server, while (Yang et al., 2024a) focuses on one-shot semi-supervised FL. However, these approaches are either inefficient or pose risks of client information leakage. In contrast, FedBiP functions as an OSFL algorithm, offering enhanced efficiency and robust privacy-preserving capabilities.
@@ 2.2 Diffusion Models For Image Synthesis
Diffusion models (Ho et al., 2020), especially Latent Diffusion Model (LDM) (Rombach et al., 2022), have attracted significant attention due to their capability to generate high-resolution natural images. They have demonstrated effectiveness in various applications, including image stylization (Guo et al., 2023; Meng et al., 2021; Kawar et al., 2023) and training data generation (Yuan et al., 2023; Sarıyıldız et al., 2023; Azizi et al., 2023). We refer readers to (Croitoru et al., 2023; Yang et al., 2023b) for a comprehensive overview of recent progress on diffusion models. Pretrained LDM has been adopted to address client data scarcity in OSFL (Zhang et al., 2023; Yang et al., 2024b). However, these methods often overlook the feature distribution shift between the LDM pretraining dataset and the clients’ local data. This challenge is particularly pronounced in complex domains such as medical and satellite imaging. To address this issue, we propose FedBiP, which personalizes the pretrained LDM to synthesize data that is aligned with the clients’ data distributions.
@@ 3.1 Heterogeneous One-shot Federated Learning
In this section, we introduce our problem setting, i.e., heterogeneous One-Shot Federated Learning (OSFL). Following (Zhang et al., 2023), we focus on image classification tasks with the goal of optimizing a $C$-way classification model $\phi$ utilizing the client local data, where $C\in\mathbb{N}$ denotes the number of categories. We assume there are $K\in\mathbb{N}$ clients joining the collaborative training. Each client $k$ owns its private dataset $D^{k}$ containing $N^{k}\in\mathbb{N}$ (image, label) pairs: $\{x^{k}_{i},y^{k}_{i}\}_{i=1}^{N^{k}}$. Only one-shot data upload from the clients to the central server is allowed.
As described in (Kairouz et al., 2021), OSFL with data heterogeneity is characterized by distribution shifts in local datasets: $P_{\mathcal{XY}}^{k_{1}}\neq P_{\mathcal{XY}}^{k_{2}}$ with $k_{1}\neq k_{2}$, where $P_{\mathcal{XY}}^{k}$ defines the joint distribution of input space $\mathcal{X}$ and label space $\mathcal{Y}$ on $D^{k}$. Data heterogeneity can be decomposed into two types: (1) label space heterogeneity, where $P_{\mathcal{Y}}$ varies across clients, while $P_{\mathcal{X|Y}}$ remains the same, and (2) feature space heterogeneity, where $P_{\mathcal{X}}$ or $P_{\mathcal{X|Y}}$ varies across clients, while $P_{\mathcal{Y|X}}$ or $P_{\mathcal{Y}}$ remains the same.
@@ 3.2 Latent Diffusion Model Pipeline
In this section, we introduce the training and inference pipelines for Latent Diffusion Model (LDM). We provide a schematic illustration in Figure 2. Given an image $x\in\mathbb{R}^{H\times W\times 3}$, the encoder $\mathcal{E}$ encodes $x$ into a latent representation $z(0)=\mathcal{E}(x)$, where $z(0)\in\mathbb{R}^{h\times w\times c}$. Besides, the decoder $\mathcal{D}$ reconstructs the image from the latent, giving $\tilde{x}=\mathcal{D}(z(0))=\mathcal{D}(\mathcal{E}(x))$. The forward diffusion and denoising processes occur in the latent representation space, as described below.
@@
Figure 2: Schematic illustration of the Latent Diffusion Model pipeline with textual prompt conditioning.
4 Methodology
4.1 Motivational Case Study
To substantiate the necessity of the proposed method, we present an empirical analysis to address the following research question: Can pretrained Latent Diffusion Model (LDM) generate images that are infrequently represented in the pretraining dataset using solely textual conditioning? Specifically, we adopt two datasets, namely DomainNet (Peng et al., 2019) and DermaMNIST (Yang et al., 2023a), which contain images indicating different styles and images from challenging medical domains, respectively. We prompt LDM with ”A quickdraw style of an airplane.” to generate airplane images in quickdraw style for DomainNet dataset, and ”A dermatoscopic image of a dermatofibroma, a type of pigmented skin lesions.” for DermaMNIST. We synthesize 100 images for each setting and adopt a pretrained ResNet-18 (He et al., 2016) to acquire the feature embeddings of real and synthetic images. Finally, we visualize them using UMAP (McInnes et al., 2018).
As shown in Figure 1, we observe markedly different visual characteristics between synthetic and real images. Specifically, for DomainNet, there exist significant discrepancies between the ”quickdraw” concept demonstration in the original dataset and the pretrained LDM. For DermaMNIST, the pretrained LDM is only able to perceive the general concepts of ”dermatoscopic” and ”skin lesion”, failing to capture category-specific information. This further highlights the difficulties in reproducing medical domain data via LDM. Additionally, there is a substantial gap in the extracted feature embeddings between real and synthetic images. Most importantly, despite the high visual quality of the synthetic images, they may not contribute to the final performance of the classification model. As demonstrated by our experimental results (Table 5.3), directly applying such prompts to generate images for server-side training sometimes yields worse results than baseline methods. Therefore, it is essential to design a more sophisticated method to effectively personalize the pretrained LDM to the specific domains of client local datasets. These observations motivate our proposed method FedBiP, which mitigates the distribution shifts between pretrained LDM and the client local data. We introduce FedBiP in the following.
4.2 Proposed Method
4.2.1 Instance-level Personalization
4.2.2 Concept-level Personalization
where timestep $t$ is sampled from $\text{Uniform}(\{1,...,T\})$.
$L_{cls}=L_{CE}(\phi(\tilde{x}),y).$ (6)
@@ 4.2.2 Concept-level Personalization
5 Experiments and Analyses
5.1 Benchmark Experiments
Datasets Description: We adapt three common image classification benchmarks with feature distribution shift for our OSFL setting: (1) DomainNet (Peng et al., 2019), which contains six domains: Clipart (C), Infograph (I), Painting (P), Quickdraw (Q), Real (R), and Sketch (S). We select 10 categories following (Zhang et al., 2023). (2) PACS (Li et al., 2017), which includes images that belong to 7 classes from four domains: Art (A), Cartoon (C), Photo (P), and Sketch (S). (3) OfficeHome (Venkateswara et al., 2017) comprises images of daily objects from four domains: Art (A), Clipart (C), Product (P), and Real (R). Each client is assigned a specific domain. To simulate local data scarcity described in previous sections, we adopt 16-shot per class (8-shot for OfficeHome) for each client, following previous works (Li et al., 2021; Chen et al., 2023).
Baseline Methods: We compare FedBiP with several baseline methods, including FedAvg and Central, i.e., aggregating the training data from all clients. We note that Central is an oracle method as it infringes on privacy requirements, while FedAvg requires multi-round communication and is not applicable to OSFL. Besides, we validate concurrent generation-based methods for OSFL: (1) FedD3 (Song et al., 2023), where distilled instances from the clients are uploaded. (2) DENSE (Zhang et al., 2022), where client local models are uploaded and distilled into one model using synthetic images. (3) FedDEO (Yang et al., 2024b), where the optimized category descriptions are uploaded and guide pretrained diffusion models. (4) FGL (Zhang et al., 2023), where captions of client local images, extracted by BLIP-2 (Li et al., 2023), are uploaded and guide pretrained LDM.
Implementation Details: We adopt the HuggingFace open-sourced ”CompVis/stable-diffusion-v1-4” as the pretrained Latent Diffusion Model, and use ResNet-18 pretrained on ImageNet (Deng et al., 2009) as the initialization for the classification model. We investigate three variants of FedBiP, namely ”S”, ”M”, and ”L”, which corresponds to generating 2×2$\times$2 ×, 5×5$\times$5 ×, 10×10$\times$10 × the number of images in the original client local dataset, respectively. Note that synthesizing more images does not affect the client’s local optimization costs. We optimize the concept vectors for 50 epochs at each client. For FGL, 3500 samples per class per domain are generated. For FedDEO, the total number of synthetic images is identical to FedBiP-L for a fair comparison. Further details about training hyperparameters are provided in the Appendix.
@@ 5.1 Benchmark Experiments
5.2 Validation on Medical and Satellite Image Datasets

### v2

@@ Abstract
One-Shot Federated Learning (OSFL), a special decentralized machine learning paradigm, has recently gained significant attention. OSFL requires only a single round of client data or model upload, which reduces communication costs and mitigates privacy threats compared to traditional FL. Despite these promising prospects, existing methods face challenges due to client data heterogeneity and limited data quantity when applied to real-world OSFL systems. Recently, Latent Diffusion Models (LDM) have shown remarkable advancements in synthesizing high-quality images through pretraining on large-scale datasets, thereby presenting a potential solution to overcome these issues. However, directly applying pretrained LDM to heterogeneous OSFL results in significant distribution shifts in synthetic data, leading to performance degradation in classification models trained on such data. This issue is particularly pronounced in rare domains, such as medical imaging, which are underrepresented in LDM’s pretraining data. To address this challenge, we propose Federated Bi-Level Personalization (FedBiP), which personalizes the pretrained LDM at both instance-level and concept-level. Hereby, FedBiP synthesizes images following the client’s local data distribution without compromising the privacy regulations. FedBiP is also the first approach to simultaneously address feature space heterogeneity and client data scarcity in OSFL. Our method is validated through extensive experiments on three OSFL benchmarks with feature space heterogeneity, as well as on challenging medical and satellite image datasets with label heterogeneity. The results demonstrate the effectiveness of FedBiP, which substantially outperforms other OSFL methods. Our code is available at https://github.com/HaokunChen245/FedBiP.
@@ 1 Introduction
Federated Learning (FL) [29] is a decentralized machine learning paradigm, in which multiple clients collaboratively train neural networks without centralizing their local data. However, traditional FL frameworks require frequent communication between a server and clients to transmit model weights, which would lead to significant communication overheads [16]. Additionally, such frequent communication increases system susceptibility to privacy threats, as transmitted data can be intercepted by attackers who may then execute membership inference attacks [27]. In contrast, a special variant of FL, One-Shot Federated Learning (OSFL) [10], serves as a promising solution. OSFL requires only single-round server-client communication, thereby enhancing communication efficiency and significantly reducing the risk of interception by malicious attackers. Therefore, we focus on OSFL given its promising properties.
@@
Figure 1: Feature map visualization of client images (real), synthetic images by prompted pretrained LDM (pretrained), and our method (FedBiP) on two datasets. FedBiP mitigates the strong distribution shifts between pretrained LDM and client local data.
Despite these promising prospects, existing methods for OSFL face significant challenges when applied to real-world scenarios. Previous works [10, 23] require additional public datasets, presenting challenges in privacy-critical domains such as medical data [26], where acquiring data that conforms to client-specific distributions is often impractical. Alternatively, they can involve the transmission of entire model weights [50] or local training data [53], which are inefficient and increase the risk of privacy leakage. Moreover, these approaches overlook the issue of feature space heterogeneity, wherein the data features across different clients exhibit non-identically distributed properties. This presents an important and prevalent challenge as emphasized in [24, 3]. Another vital challenge in (One-Shot) FL is the limited quantity of data available from clients [29]. This problem is particularly notable in specialized domains, such as medical or satellite imaging [37] where data collection is time-consuming and costly.
@@
Data augmentation constitutes a promising strategy to address these challenges in traditional FL [54, 25] by optimizing an auxiliary generative model. However, its reliance on multiple communication rounds makes it unsuitable for OSFL. Recently, diffusion models [14], particularly Latent Diffusion Model (LDM) [33], have gained significant attention due to their capability to synthesize high-quality images after being pretrained on large-scale datasets. They are proven effective in various tasks, including training data augmentation [48, 1] and addressing feature shift problems [31, 8] under centralized settings. However, directly applying a pretrained LDM for specialized domains presents challenges. As demonstrated in Figure 1, there is a noticeable feature distributional shift and visual discrepancy between real and synthetic data. This mismatch could lead to performance degradation when incorporating such synthetic data into training, especially in heterogeneous OSFL, where each client possesses data with varying distributions.
@@
Therefore, in this paper, we propose Federated Bi-Level Personalization (FedBiP), a framework designed to adapt pretrained LDM for synthesizing high-quality training data that adheres to client-specific data distributions in OSFL. FedBiP incorporates personalization of the pretrained LDM at both instance and concept levels. Specifically, instance-level personalization focuses on adapting the pretrained LDM to generate high-fidelity samples that closely align with each client’s local data while preserving data privacy. Concurrently, concept-level personalization integrates category and domain-specific concepts from different clients to enhance data generation diversity at the central server. This bi-level personalization approach improves the classification models trained on the synthesized data. Our contributions can be summarized as follows:
@@ 2.1 One-Shot Federated Learning
A variety of efforts have been made to address One-Shot Federated Learning (OSFL), primarily from two complementary perspectives: one focuses on model aggregation through techniques such as model prediction averaging [10], majority voting [23], conformal prediction method [15], loss surface adaptation [39], or Bayesian methods [49, 4, 12]. These approaches may not fully exploit the underlying knowledge across different client data distributions. Another aims to transmit training data instead of model weights: data distribution [19, 2, 36], Generative Adversarial Networks (GANs) [9, 50, 18, 17, 6], or distilled dataset [53, 38] are optimized and transmitted to the central server for subsequent model training. Given the success of diffusion models [33], [51, 45] suggests transmitting image captions to reproduce training data at the server, while [44] focuses on one-shot semi-supervised FL. However, these approaches are either inefficient or pose risks of client information leakage. In contrast, FedBiP functions as an OSFL algorithm, offering enhanced efficiency and robust privacy-preserving capabilities.
@@ 2.2 Diffusion Models For Image Synthesis
Diffusion models [14], especially Latent Diffusion Model (LDM) [33], have attracted significant attention due to their capability to generate high-resolution natural images. They have demonstrated effectiveness in various applications, including image stylization [11, 30, 20] and training data generation [48, 35, 1]. We refer readers to [5, 43] for a comprehensive overview of recent progress on diffusion models. Pretrained LDM has been adopted to address client data scarcity in OSFL [51, 45]. However, these methods often overlook the feature distribution shift between the LDM pretraining dataset and the clients’ local data. This challenge is particularly pronounced in complex domains such as medical and satellite imaging. To address this issue, we propose FedBiP, which personalizes the pretrained LDM to synthesize data aligned with the client data distributions.
@@ 3.1 Heterogeneous One-shot Federated Learning
In this section, we introduce our problem setting, i.e., heterogeneous One-Shot Federated Learning (OSFL). Following [51], we focus on image classification tasks with the goal of optimizing a $C$-way classification model $\phi$ utilizing the client local data, where $C\in\mathbb{N}$ denotes the number of categories. We assume there are $K\in\mathbb{N}$ clients joining the collaborative training. Each client $k$ owns its private dataset $D^{k}$ containing $N^{k}\in\mathbb{N}$ (image, label) pairs: $\{x^{k}_{i},y^{k}_{i}\}_{i=1}^{N^{k}}$. Only one-shot data upload from the clients to the central server is allowed.
As described in [16], OSFL with data heterogeneity is characterized by distribution shifts in local datasets: $P_{\mathcal{XY}}^{k_{1}}\neq P_{\mathcal{XY}}^{k_{2}}$ with $k_{1}\neq k_{2}$, where $P_{\mathcal{XY}}^{k}$ defines the joint distribution of input space $\mathcal{X}$ and label space $\mathcal{Y}$ on $D^{k}$. Data heterogeneity can be decomposed into two types: (1) label space heterogeneity, where $P_{\mathcal{Y}}$ varies across clients, while $P_{\mathcal{X|Y}}$ remains the same, and (2) feature space heterogeneity, where $P_{\mathcal{X}}$ or $P_{\mathcal{X|Y}}$ varies across clients, while $P_{\mathcal{Y|X}}$ or $P_{\mathcal{Y}}$ remains the same.
@@ 3.2 Latent Diffusion Model Pipeline
In this section, we introduce the training and inference pipelines for Latent Diffusion Model (LDM). Given an image $x\in\mathbb{R}^{H\times W\times 3}$, the encoder $\mathcal{E}$ encodes $x$ into a latent representation $z(0)=\mathcal{E}(x)$, where $z(0)\in\mathbb{R}^{h\times w\times c}$. Besides, the decoder $\mathcal{D}$ reconstructs the image from the latent, giving $\tilde{x}=\mathcal{D}(z(0))=\mathcal{D}(\mathcal{E}(x))$. The forward diffusion and denoising processes occur in the latent representation space, as described below.
@@
4 Methodology
4.1 Motivational Case Study
To substantiate the necessity of the proposed method, we present an empirical analysis to address the following research question: Can pretrained Latent Diffusion Model (LDM) generate images that are infrequently represented in the pretraining dataset using solely textual conditioning? Specifically, we adopt two datasets, namely DomainNet [32] and DermaMNIST [42], which contain images indicating different styles and images from challenging medical domains, respectively. We prompt LDM with ”A quickdraw style of an airplane.” to generate airplane images in quickdraw style for DomainNet dataset, and ”A dermatoscopic image of a dermatofibroma, a type of pigmented skin lesions.” for DermaMNIST. We synthesize 100 images for each setting and adopt a pretrained ResNet-18 [13] to acquire the feature embeddings of real and synthetic images. Finally, we visualize them using UMAP [28].
As shown in Figure 1, we observe markedly different visual characteristics between synthetic and real images. Specifically, for DomainNet, there exist significant discrepancies between the ”quickdraw” concept demonstration in the original dataset and the pretrained LDM. For DermaMNIST, the pretrained LDM is only able to perceive the general concepts of ”dermatoscopic” and ”skin lesion”, failing to capture category-specific information. This further highlights the difficulties in reproducing medical domain data via LDM. Additionally, there is a substantial gap in the extracted feature embeddings between real and synthetic images. Most importantly, despite the high visual quality of the synthetic images, they may not contribute to the final performance of the classification model. As demonstrated by our experimental results (Table 3), directly applying such prompts to generate images for server-side training sometimes yields worse results than baseline methods. Therefore, it is essential to design a more sophisticated method to effectively personalize the pretrained LDM to the specific domains of client local datasets. These observations motivate our proposed method FedBiP, which mitigates the distribution shifts between pretrained LDM and client local data. We introduce FedBiP in the following.
4.2 Proposed Method
4.2.1 Instance-level Personalization
@@
4.2.2 Concept-level Personalization
where timestep $t$ is sampled from $\text{Uniform}(\{1,...,T\})$.
$L_{cls}=L_{CE}(\phi(\tilde{x}),y).$ (6)
5 Experiments and Analyses
5.1 Benchmark Experiments
Datasets Description: We adapt three common image classification benchmarks with feature distribution shift for our OSFL setting: (1) DomainNet [32], which contains six domains: Clipart (C), Infograph (I), Painting (P), Quickdraw (Q), Real (R), and Sketch (S). We select 10 categories following [51]. (2) PACS [21], which includes images that belong to 7 classes from four domains: Art (A), Cartoon (C), Photo (P), and Sketch (S). (3) OfficeHome [41] comprises images of daily objects from four domains: Art (A), Clipart (C), Product (P), and Real (R). Each client is assigned a specific domain. To simulate local data scarcity described in previous sections, we adopt 16-shot per class (8-shot for OfficeHome) for each client, following [24, 3].
Baseline Methods: We compare FedBiP with several baseline methods, including FedAvg and Central, i.e., aggregating the training data from all clients. We note that Central is an oracle method as it infringes on privacy requirements, while FedAvg requires multi-round communication and is not applicable to OSFL. Besides, we validate concurrent generation-based methods for OSFL: (1) FedD3 [38], where distilled instances from the clients are uploaded. (2) DENSE [50], where client local models are uploaded and distilled into one model using synthetic images. (3) FedDEO [45], where the optimized category descriptions are uploaded and guide pretrained diffusion models. (4) FGL [51], where captions of client local images, extracted by BLIP-2 [22], are uploaded and guide pretrained LDM.
Implementation Details: We adopt the HuggingFace open-sourced ”CompVis/stable-diffusion-v1-4” as the pretrained Latent Diffusion Model, and use ResNet-18 pretrained on ImageNet [7] as the initialization for the classification model. We investigate three variants of FedBiP, namely ”S”, ”M”, and ”L”, which corresponds to generating 2×2$\times$2 ×, 5×5$\times$5 ×, 10×10$\times$10 × the number of images in the original client local dataset, respectively. Note that synthesizing more images does not affect the client’s local optimization costs. We optimize the concept vectors for 50 epochs at each client. For FGL, 3500 samples per class per domain are generated. For FedDEO, the total number of synthetic images is identical to FedBiP-L for a fair comparison. Further details about training hyperparameters are provided in Appendix.